A 15.46 g sample of a compound contains 9.87 g of iron, Fe, 1.82 g of phosphorus, P, and oxygen, O. Calculate the empirical formula for the compound.
Answer #1

15.46 g sample of a compound contains 9.87 g of iron, Fe, 1.82 g of phosphorus, P, and oxygen

Atomic mass of Fe = 56 gm/mol

Atomic mass of O = 16 gm/mol

Atomic mass of P= 31 gm/mol

mass of Fe = 9.87 gm

mass of  P = 1.82 gm

mass of O = 15.46 -(9.87 + 1.82) = 15.46 - 11.69 = 3.77 gm

number of moles = mass /atomic mass

(moles )O = 3.77 gm / 16gm/mol= 0.2356 moles

(moles )Fe = 9.87 gm / 56gm/mol= 0.17625 moles

(moles )P = 1.82gm / 31 gm/mol = 0.05871 moles

Fe O P

Moles: 0.17625 0.2356 0.05871  

Dividing with the smallest number i.e.- 0.05871 we have,

Moles:   3 4 1

Thus, the empirical formula is Fe3O4P